# Statistical Response of ENSO Complexity to Initial Value and Model Parameter Perturbations

### v1

@@ Abstract
The purpose of this work is to better understand how the El Niño-Southern Oscillation (ENSO) responds to changes in its initial state and internal dynamics or external forcings. A statistical quantification of this response allows for the comprehension of the triggering conditions and the effect of climate change in the occurrence frequency and strength of each type of ENSO event. Such a study also allows to detect the most dangerous perturbation directions, which has practical significance in guiding anthropogenic activities. The approach used to study the response in this work is through the framework of information theory, which allows for an unbiased and robust assessment of the statistical response that is not affected by the turbulent dynamics of the system.
@@ 1 Introduction
Due to the strong connections with other climate variabilities, it is of practical importance to compute the response of each ENSO event to various perturbations of the climate system. On the one hand, predicting the corresponding spatiotemporal patterns of an ENSO event due to the perturbed initial state helps understand its precursors and analyze the triggering conditions of the event. Such a study is vital in advancing an improved understanding of ENSO predictability. It also facilitates discovering the formation mechanisms of extreme El Niños and thus provides possible guidance to prevent or cope with the associated natural hazards. On the other hand, a perturbation of either the internal dynamics or the external forcing will also lead to a change in the resulting ENSO spatiotemporal patterns. Such a perturbation can be caused by potential climate change and may result in regime switching. The perturbed system may bring about a different occurrence frequency for each type of ENSO event. It may also increase the strength of extreme events and enhance the probability of generating multi-year events. With an appropriate climate model, a natural way to study these two types of perturbation problems is to compute the model response to the perturbed initial conditions and model parameters, respectively. Particularly, detecting the most dangerous perturbation directions, namely the fastest-growing perturbations, has practical significance in guiding anthropogenic activities.
@@ 2.2 Statistical response and the associated information measurement
Given a perturbation, the strength of the corresponding statistical response is computed from the relative entropy in (2). However, unless the distributions have desirable features, numerical integration is needed to calculate the relative entropy, which is a computationally challenging issue. Furthermore, in the situation of seeking the most dangerous perturbation directions, an exhaustive search of the entire state space of $\mathbf{x}_{t}$ is needed based on the direct definition of the relative entropy in (2). This becomes computationally prohibitive when $\mathbf{x}_{t}$ is high dimensional (Robert et al., 2010; Kuo and Sloan, 2005). The following subsection aims to provide alternative ways to accelerate the calculations.
@@ 2.3.2 Leading-order approximation via Fisher information
Recall that $\bm{\delta}$ is the perturbation vector, which can be the perturbation of a subset of the state variables for the initial values or a few selected parameters. It is assumed that the possible range for the perturbation is within physical meanings and that it is further standardized as for $\bm{\delta}=\mathbf{0}$ to correspond to the unperturbed climatological system. Since the perturbation is usually small, the perturbed PDF can be written as a function of $\bm{\delta}$. Applying a Taylor expansion of $p^{\delta}(\mathbf{x}_{t})$ with respect to $\bm{\delta}$ in computing the relative entropy under the tacit assumption that the PDF is differentiable with respect to the perturbation $\bm{\delta}$ (Majda et al., 2009; Majda and Wang, 2010; Hairer and Majda, 2010), yields the following leading-order approximation of the response to the perturbation (Majda and Chen, 2018; Majda and Gershgorin, 2011),
@@
where the first term on the right-hand side of (5) is the quadratic form in $\bm{\delta}$ given by the Fisher information (Williams, 2001; Cover, 1999),
@@ 2.3.3 Fisher information with coarse-grained statistical measurements
In many situations, the observed climatology data are used to compute unperturbed statistics. Yet, it is worth noting that, due to the limited amount of data and the possible measurement noise, these data can usually estimate the first few moments accurately, but the higher-order moments are very sensitive to the small noise. Therefore, instead of calculating the exact PDF, the measured leading few moments are typically used to reconstruct the least biased PDF using the so-called maximum entropy principle (Majda and Wang, 2006; Bajkova, 1992). It is then used as an approximation of $p(\mathbf{x}_{t})$ in (6) to compute the most dangerous perturbation direction. See Majda et al. (2005), Majda and Wang (2006), and Majda and Gershgorin (2010) for more details.
Denote by $\mathbf{E}_{L}(\mathbf{x}_{t})=(E_{1}(\mathbf{x}_{t}),\ldots,E_{L}(\mathbf{x}_{t}))$ the $L$ statistical quantities from the observational measurements or model simulation, for example, the mean, the covariance and up to the $L$-th moment. The least biased PDF obtained from the maximum entropy principle is denoted by $p^{{\delta}}_{L}(\mathbf{x}_{t})$. Further denote by $\overline{\mathbf{E}}_{L}=(\overline{E}_{1},\ldots,\overline{E}_{L})$, where each component $\overline{{E}}_{l}$ is given by $\overline{{E}}_{l}=\int_{\mathbf{x}_{t}}{E}_{l}(\mathbf{x}_{t})p^{{\delta}}_{L}(\mathbf{x}_{t})\mathrm{d}\mathbf{x}_{t}$. The notation ‘overline’ represents the statistical average with respect to the perturbed PDF, and the resulting $\overline{{E}}_{l}$ is a number that depends on the perturbed dynamics. Then the quadratic form of $\bm{\delta}\cdot I(p(\mathbf{x}_{t}))\bm{\delta}$ in (5) can be approximated by (Majda and Gershgorin, 2010)
$\bm{\delta}\cdot I(p_{L}(\mathbf{x}_{t}))\bm{\delta}=\bm{\delta}\cdot\left(\nabla_{\bm{\delta}}\overline{\mathbf{E}}_{L}\mathcal{C}^{-1}\nabla_{\bm{\delta}}\overline{\mathbf{E}}_{L}^{\mathtt{\,T}}\right)\bm{\delta},$ (7)
@@
and the Jacobian of the statistical average is evaluated at the unperturbed state. Note that ${E}_{l}(\mathbf{x}_{t})$ is a function of the state variable $\mathbf{x}_{t}$ while its statistical average $\overline{{E}}_{l}$ is a number. For example, if the mean and the variance are adopted as the first two components of the measurements, then ${E}_{1}(\mathbf{x}_{t})=\mathbf{x}_{t}$ and ${E}_{2}(\mathbf{x}_{t})=(\mathbf{x}_{t}-\overline{\mathbf{x}}_{t})^{2}$. Correspondingly, $\overline{E}_{1}(\mathbf{x}_{t})=\overline{\mathbf{x}}_{t}$ and $\overline{E}_{2}(\mathbf{x}_{t})=\overline{(\mathbf{x}_{t}-\overline{\mathbf{x}}_{t})^{2}}$. The associated first four entries of $\mathcal{C}$ in (8) are given by
@@
$\bm{\delta}\cdot I(p_{L}(\mathbf{x}_{t}))\bm{\delta}=\bm{\delta}\cdot\left(\nabla_{\bm{\delta}}\overline{\mathbf{E}}_{L^{\prime}}\mathcal{C}^{-1}\nabla_{\bm{\delta}}\overline{\mathbf{E}}_{L^{\prime}}^{\mathtt{\,T}}\right)\bm{\delta},$ (9)
where $\mathbf{E}_{L^{\prime}}(\mathbf{x}_{t})=(E_{1}(\mathbf{x}_{t}),\ldots,E_{L^{\prime}}(\mathbf{x}_{t}),0,\ldots,0))$. That is, the first $L^{\prime}$ entries of $\mathbf{E}_{L^{\prime}}(\mathbf{x}_{t})$ are the same as those in $\mathbf{E}_{L}(\mathbf{x}_{t})$ but the remaining entries are zero. The compressed quadratic form in (9) is relevant in determining the important practical information regarding whether, for example, changes in the mean climate statistics alone determine the most sensitive directions of climate change.
@@ 2.4 Summary
Note that, although the quad form w/ mean and variance utilizes the Gaussian statistics, it differs from applying the Gaussian approximation (4) to the exact formula. The quadratic form via the Fisher information (6) is already an approximation in computing the statistical response since the Taylor expansion of the relative entropy in (5) is utilized. Building upon this, the first two moments are adopted to replace the full PDF as a second approximation. Nevertheless, it is worth highlighting that the methods exploiting the quadratic form facilitate determining the most sensitive (i.e., dangerous) perturbation direction by finding the unit eigenvector corresponding to the largest eigenvalue of the matrix $I(p(\mathbf{x}_{t}))$ or $I(p_{L}(\mathbf{x}_{t}))$, hereafter named the maximal eigenvector.
@@
The statistical quantities in this work are computed from an ensemble simulation. It is based on a Monte Carlo simulation with $3000$ ensemble members. Such an ensemble size is large enough to reproduce the strong non-Gaussian climatology PDFs for the 6-dimensional conceptual model described in Section 3. The computational cost of using $3000$ ensemble members remains low for such a conceptual model. Therefore, it provides an accurate reference solution of the statistical response using the exact formula. It further allows us to compare the result with those using the Gaussian approximations and the quad form w/ mean and variance. Note that a small ensemble size is usually sufficient to reach reasonably accurate results when applying the Gaussian approximation or the Gaussian statistics in the quadratic form. Therefore, these approximate methods can compute the statistical response using more sophisticated and higher-dimensional operational systems. The unperturbed initial value is always given by a Gaussian distribution, centered at the observational values and equipped with a tiny variance of $10^{-4}$ along each direction. Adding such a small uncertainty to the initial value facilitates the numerical calculation of the relative entropy within a very short lead time. Other advanced techniques, such as the fluctuation-dissipation theorem (Majda et al., 2005; Kubo, 1966), can be embedded into the above methods to compute the statistical response more efficiently. These methods are helpful for more complicated systems but are not adopted here since the direct Monte Carlo simulation with $3000$ ensemble members is sufficient for the 6-dimensional conceptual model. Finally, when calculating $\nabla_{\bm{\delta}}p(\mathbf{x}_{t})$ in (6), where $p(\mathbf{x}_{t})$ is the unperturbed distribution, the derivative is approximated by a second-order accurate centered finite difference. This is achieved by adding a small numerical perturbation $\bm{\epsilon}$ with $\|\bm{\epsilon}\|\ll\|\bm{\delta}\|$ to the initial conditions or model parameters, and then computing the associated PDF at time $t$. Likewise, compute another PDF at time $t$ corresponding to the perturbation $-\bm{\epsilon}$. These two PDFs are then used to approximate the derivative numerically as
@@ 3.1 Observational data sets
The monthly ocean temperature and current data are from the GODAS reanalysis dataset (Behringer and Xue, 2004). The thermocline depth along the equatorial Pacific is approximated from the potential temperature as the depth of the $20^{\circ}$ isotherm contour. The analysis period is 36 years, from the start of 1982 until the end of 2017. Anomalies presented in this study are calculated by removing the monthly mean climatology of the whole period. In this work, the Niño 4 ($T_{C}$) and Niño 3 ($T_{E}$) indices are the averages of the SSTa over the CP ($160^{\circ}$-$150^{\circ}$, $5^{\circ}$-$5^{\circ}$) and EP ($150^{\circ}$-$90^{\circ}$, $5^{\circ}$-$5^{\circ}$) regions, respectively. The $h_{W}$ index is the mean thermocline depth anomaly over the western Pacific region ($120^{\circ}$-$180^{\circ}$, $5^{\circ}$-$5^{\circ}$), while the $u$ index is the mean mixed-layer zonal current in the CP region.
@@ 3.3 The multiscale stochastic dynamical model for the ENSO complexity
The dynamical core of the model is a deterministic three-region interannual linear model with zonal advective feedback (Fang and Mu, 2018). It extends on the classical recharge oscillator model (Jin, 1997) and implements the air-sea interactions over the entire WP, CP, and EP regions. It also incorporates the ocean content discharge and recharge process controlling the occurrence of El Niño and La Niña events via the thermal layer and the ocean zonal advection. In the model, $T_{C}$ and $T_{E}$ are the SSTa in the CP and EP regions, respectively, while $u$ is the mean zonal current anomaly in the CP region and $h_{W}$ is the mean thermocline depth anomaly in the WP region. In addition to these interannual variabilities, two processes, describing the intraseasonal wind bursts $\tau$ and the decadal variability in the background Walker circulation $I$, are further incorporated into the model. The intraseasonal variability $\tau$ accounts for several important atmospheric ENSO triggers, such as the westerly wind bursts (WWBs), the easterly wind bursts (EWBs), and the convective envelope of the Madden-Julian Oscillation (MJO) (Chen et al., 2015; Hu and Fedorov, 2016; Puy et al., 2016; Vecchi et al., 2006). Its strength is given by a state-dependent (multiplicative) noise that depends on the SSTa (Jin et al., 2007; Bianucci et al., 2018; Chen and Zhang, 2023), where a warmer SSTa leads to stronger wind burst activities. As for the decadal variability in the state variables through $I$, it stems from the observation that since 1870, through several detailed El Niño-type classification methods, the EP and CP El Niño events are alternatively prevalent every 10 to 20 years (Yu and Kim, 2013; Dieppois et al., 2021). This oscillation between EP-dominant and CP-dominant regimes indicates that the decadal variability plays an important role in the underlying dynamics, which is parameterized through a simple linear stochastic differential equation with multiplicative noise (10f), with no explicit dependence on the state variables in the faster time scales (Yang et al., 2021). They, together with additional small Gaussian white noise $\sigma_{u}\dot{W}_{u}$, $\sigma_{h}\dot{W}_{h}$, $\sigma_{C}\dot{W}_{C}$, $\sigma_{E}\dot{W}_{E}$, characterize the irregularity and multiscale features of the ENSO complexity (Timmermann et al., 2018; Fang and Xie, 2020).
@@
In the following, the non-dimensional form of the model in (3.3) is utilized to compute the responses. In the non-dimensional form, the six state variables have comparable maximum amplitudes. This means the strength of the components in the 6-dimensional column vector describing the eigenvector of the quadratic form can be used to intuitively tell the most sensitive direction. Nevertheless, the results for the evolution of the statistics with or without the perturbation are shown in the dimensional form that facilitates the physical explanations.

### v2

# Statistical Response of ENSO Complexity to Initial Condition and Model Parameter Perturbations

@@ Abstract
The purpose of this work is to better understand how the El Niño-Southern Oscillation (ENSO) responds to changes in its initial state and internal dynamics or external forcings. A statistical quantification of this response allows for the comprehension of the triggering conditions and the effect of climate change in the occurrence frequency and strength of each type of ENSO event. Such a study also allows to detect the most sensitive perturbation directions, which has practical significance in guiding anthropogenic activities. The approach used to study the response in this work is through the framework of information theory, which allows for an unbiased and robust assessment of the statistical response that is not affected by the turbulent dynamics of the system.
@@ 1 Introduction
Due to the strong connections with other climate variabilities, it is of practical importance to compute the response of each ENSO event to various perturbations of the climate system. On the one hand, predicting the corresponding spatiotemporal patterns of an ENSO event due to the perturbed initial state helps understand its precursors and analyze the triggering conditions of the event. For example, calculating the response of the SST field due to the perturbed wind stress field is one of the most essential practical topics. Such a study is vital in advancing an improved understanding of ENSO predictability. It also facilitates discovering the formation mechanisms of extreme El Niños and thus provides possible guidance to prevent or cope with the associated natural hazards. On the other hand, a perturbation of either the internal dynamics or the external forcing will also lead to a change in the resulting ENSO spatiotemporal patterns. Such a perturbation can be caused by potential climate change and may result in regime switching. The perturbed system may bring about a different occurrence frequency for each type of ENSO event. It may also increase the strength of extreme events and enhance the probability of generating multi-year events. With an appropriate climate model, a natural way to study these two types of perturbation problems is to compute the model response to the perturbed initial conditions and model parameters, respectively. Particularly, detecting the most sensitive perturbation directions, namely the fastest-growing perturbations, has practical significance in guiding anthropogenic activities.
@@ 2.2 Statistical response and the associated information measurement
Given a perturbation, the strength of the corresponding statistical response is computed from the relative entropy in (2). However, unless the distributions have desirable features, numerical integration is needed to calculate the relative entropy, which is a computationally challenging issue. Furthermore, in the situation of seeking the most sensitive perturbation directions, an exhaustive search of the entire state space of $\mathbf{x}_{t}$ is needed based on the direct definition of the relative entropy in (2). This becomes computationally prohibitive when $\mathbf{x}_{t}$ is high dimensional (Robert et al., 2010; Kuo and Sloan, 2005). The following subsection aims to provide alternative ways to accelerate the calculations.
@@ 2.3.2 Leading-order approximation via Fisher information
Recall that $\bm{\delta}$ is the perturbation vector, which can be the perturbation of a subset of the state variables for the initial values or a few selected parameters, and assumed to be an $N$ dimensional vector. It is assumed that the possible range for the perturbation is within physical meanings and that it is further standardized as for $\bm{\delta}=\mathbf{0}$ to correspond to the unperturbed climatological system. Since the perturbation is usually small, the perturbed PDF can be written as a function of $\bm{\delta}$. Applying a Taylor expansion of $p^{\delta}(\mathbf{x}_{t})$ with respect to $\bm{\delta}$ in computing the relative entropy under the tacit assumption that the PDF is differentiable with respect to the perturbation $\bm{\delta}$ (Majda et al., 2009; Majda and Wang, 2010; Hairer and Majda, 2010), yields the following leading-order approximation of the response to the perturbation (Majda and Chen, 2018; Majda and Gershgorin, 2011),
@@
where the first term on the right-hand side of (5) is the quadratic form in $\bm{\delta}$ defined by the $N\times N$ Fisher information matrix (Williams, 2001; Cover, 1999),
@@ 2.3.3 Fisher information with coarse-grained statistical measurements
In many situations, the observed climatology data are used to compute unperturbed statistics. Yet, it is worth noting that, due to the limited amount of data and the possible measurement noise, these data can usually estimate the first few moments accurately, but the higher-order moments are very sensitive to the small noise. Therefore, instead of calculating the exact PDF, the measured leading few moments are typically used to reconstruct the least biased PDF using the so-called maximum entropy principle (Majda and Wang, 2006; Bajkova, 1992). It is then used as an approximation of $p(\mathbf{x}_{t})$ in (6) to compute the most sensitive perturbation direction. See Majda et al. (2005), Majda and Wang (2006), and Majda and Gershgorin (2010) for more details.
Denote by $\mathbf{E}_{L}(\mathbf{x}_{t})=(E_{1}(\mathbf{x}_{t}),\ldots,E_{L}(\mathbf{x}_{t}))$ the $L$ statistical quantities from the observational measurements or model simulation, for example, the mean, the covariance and up to the $L$-th moment. The least biased PDF obtained from the maximum entropy principle is denoted by $p^{{\delta}}_{L}(\mathbf{x}_{t})$. Further denote by $\overline{\mathbf{E}}_{L}=(\overline{E}_{1},\ldots,\overline{E}_{L})^{\mathtt{\,T}}$, where each component $\overline{{E}}_{l}$ is given by $\overline{{E}}_{l}=\int_{\mathbf{x}_{t}}{E}_{l}(\mathbf{x}_{t})p^{{\delta}}_{L}(\mathbf{x}_{t})\mathrm{d}\mathbf{x}_{t}$. The notation ‘overline’ represents the statistical average with respect to the perturbed PDF, and the resulting $\overline{{E}}_{l}$ is a number that depends on the perturbed dynamics. Then the quadratic form of $\bm{\delta}\cdot I(p(\mathbf{x}_{t}))\bm{\delta}$ in (5) can be approximated by (Majda and Gershgorin, 2010)
$\bm{\delta}\cdot I(p_{L}(\mathbf{x}_{t}))\bm{\delta}=\bm{\delta}\cdot\left(\big{(}\nabla_{\bm{\delta}}\overline{\mathbf{E}}_{L}\big{)}^{\mathtt{\,T}}\mathcal{C}^{-1}\nabla_{\bm{\delta}}\overline{\mathbf{E}}_{L}\right)\bm{\delta},$ (7)
@@
and $\nabla_{\delta}\overline{\mathbf{E}}_{L}$ is the gradient of each component in $\overline{\mathbf{E}}_{L}$ with respect to the perturbation vector $\bm{\delta}$ that gives an $L\times N$ Jacobi matrix with standard element $\left(\nabla_{\delta}\overline{\mathbf{E}}_{L}\right)^{j=1,\ldots,N}_{l=1,\ldots,L}=\frac{\partial\overline{{E}}_{l}}{\partial\delta_{j}}$ and with the Jacobian of the statistical average being evaluated at the unperturbed state of $\bm{\delta}=\mathbf{0}$. Note that ${E}_{l}(\mathbf{x}_{t})$ is a function of the state variable $\mathbf{x}_{t}$ while its statistical average $\overline{{E}}_{l}$ is a number. For example, if the mean and the variance are adopted as the first two components of the measurements, then ${E}_{1}(\mathbf{x}_{t})=\mathbf{x}_{t}$ and ${E}_{2}(\mathbf{x}_{t})=(\mathbf{x}_{t}-\overline{\mathbf{x}}_{t})^{2}$. Correspondingly, $\overline{E}_{1}(\mathbf{x}_{t})=\overline{\mathbf{x}}_{t}$ and $\overline{E}_{2}(\mathbf{x}_{t})=\overline{(\mathbf{x}_{t}-\overline{\mathbf{x}}_{t})^{2}}$. The associated first four entries of $\mathcal{C}$ in (8) are given by
@@
$\bm{\delta}\cdot I(p_{L}(\mathbf{x}_{t}))\bm{\delta}=\bm{\delta}\cdot\left(\big{(}\nabla_{\bm{\delta}}\overline{\mathbf{E}}_{L^{\prime}}\big{)}^{\mathtt{\,T}}\mathcal{C}^{-1}\nabla_{\bm{\delta}}\overline{\mathbf{E}}_{L^{\prime}}\right)\bm{\delta},$ (9)
where $\mathbf{E}_{L^{\prime}}(\mathbf{x}_{t})=(E_{1}(\mathbf{x}_{t}),\ldots,E_{L^{\prime}}(\mathbf{x}_{t}),0,\ldots,0))^{\mathtt{\,T}}$. That is, the first $L^{\prime}$ entries of $\mathbf{E}_{L^{\prime}}(\mathbf{x}_{t})$ are the same as those in $\mathbf{E}_{L}(\mathbf{x}_{t})$ but the remaining entries are zero. The compressed quadratic form in (9) is relevant in determining the important practical information regarding whether, for example, changes in the mean climate statistics alone determine the most sensitive directions of climate change.
@@ 2.4 Summary
Note that, although the quad form w/ mean and variance utilizes the Gaussian statistics, it differs from applying the Gaussian approximation (4) to the exact formula. The quadratic form via the Fisher information (6) is already an approximation in computing the statistical response since the Taylor expansion of the relative entropy in (5) is utilized. Building upon this, the first two moments are adopted to replace the full PDF as a second approximation. Nevertheless, it is worth highlighting that the methods exploiting the quadratic form facilitate determining the most sensitive perturbation direction by finding the unit eigenvector corresponding to the largest eigenvalue of the matrix $I(p(\mathbf{x}_{t}))$ or $I(p_{L}(\mathbf{x}_{t}))$, hereafter named the maximal eigenvector.
@@
The statistical quantities in this work are computed from an ensemble simulation. It is based on a Monte Carlo simulation with $3000$ ensemble members. Such an ensemble size is large enough to reproduce the strong non-Gaussian climatology PDFs for the 6-dimensional conceptual model described in Section 3. The computational cost of using $3000$ ensemble members remains low for such a conceptual model. Therefore, it provides an accurate reference solution of the statistical response using the exact formula. It further allows us to compare the result with those using the Gaussian approximations and the quad form w/ mean and variance. Note that a small ensemble size is usually sufficient to reach reasonably accurate results when applying the Gaussian approximation or the Gaussian statistics in the quadratic form. Therefore, these approximate methods can compute the statistical response using more sophisticated and higher-dimensional operational systems. The unperturbed initial value is always given by a Gaussian distribution, centered at the observational values and equipped with a tiny variance of $10^{-4}$ along each direction. Adding such a small uncertainty to the initial value facilitates the numerical calculation of the relative entropy within a very short lead time. Other advanced techniques, such as the fluctuation-dissipation theorem (Majda et al., 2005; Kubo, 1966), can be embedded into the above methods to compute the statistical response more efficiently. These methods are helpful for more complicated systems but are not adopted here since the direct Monte Carlo simulation with $3000$ ensemble members is sufficient for the 6-dimensional conceptual model. Finally, when calculating $\nabla_{\bm{\delta}}p(\mathbf{x}_{t})$ in (6), where $p(\mathbf{x}_{t})$ is the unperturbed distribution, the derivative is approximated by a second-order accurate centered finite difference. This is achieved by adding a small numerical perturbation $\bm{\epsilon}$ with $\|\bm{\epsilon}\|\ll\|\bm{\delta}\|$ to the initial conditions or model parameters, and then computing the associated PDF at time $t$. Notably, the amplitude of the numerical perturbation $\bm{\epsilon}$ is required to be small to maintain the accuracy of computing the derivative and guarantee the dominant role of the actual perturbation $\bm{\delta}$ in computing the response. Likewise, we compute another PDF at time $t$ corresponding to the perturbation $-\bm{\epsilon}$. These two PDFs are then used to approximate the derivative numerically as
@@
Note that the focus of analyzing the responses to the perturbations of initial conditions and model parameters can be different. Since the initial value perturbation will have an immediate impact at short- and medium-range lead times, studying the difference in the response of each single ENSO event will be highlighted. In contrast, the short-term behavior of the system may not be affected by a sudden change in the parameter values. Therefore, parameter perturbation focuses more on long-term behavior that affects the climatology.
@@ 3.1 Observational data sets
In most of the studies in this work, the monthly ocean temperature and current data are from the GODAS reanalysis dataset (Behringer and Xue, 2004). The thermocline depth along the equatorial Pacific is approximated from the potential temperature as the depth of the $20^{\circ}$ isotherm contour. The analysis period is 36 years, from the start of 1982 until the end of 2017. Anomalies presented in this study are calculated by removing the monthly mean climatology of the whole period. In this work, the Niño 4 ($T_{C}$) and Niño 3 ($T_{E}$) indices are the averages of the SSTa over the CP ($160^{\circ}$-$150^{\circ}$, $5^{\circ}$-$5^{\circ}$) and EP ($150^{\circ}$-$90^{\circ}$, $5^{\circ}$-$5^{\circ}$) regions, respectively. The $h_{W}$ index is the mean thermocline depth anomaly over the western Pacific region ($120^{\circ}$-$180^{\circ}$, $5^{\circ}$-$5^{\circ}$), while the $u$ index is the mean mixed-layer zonal current in the CP region. In the last subsection of the results, the ERSSTv5 data (Huang et al., 2017), which has a longer period, is utilized to compute the number of each type of ENSO event.
@@ 3.3 The multiscale stochastic dynamical model for the ENSO complexity
The dynamical core of the model is a deterministic three-region interannual linear model with zonal advective feedback (Fang and Mu, 2018). It extends on the classical two-region recharge oscillator model (Jin, 1997) and implements the air-sea interactions over the entire WP, CP, and EP regions. It also incorporates the ocean content discharge and recharge process controlling the occurrence of El Niño and La Niña events via the thermal layer and the ocean zonal advection. In the model, $T_{C}$ and $T_{E}$ are the SSTa in the CP and EP regions, respectively, while $u$ is the mean zonal current anomaly in the CP region and $h_{W}$ is the mean thermocline depth anomaly in the WP region. In addition to these interannual variabilities, two processes, describing the intraseasonal wind bursts $\tau$ and the decadal variability in the background Walker circulation $I$, are further incorporated into the model. The intraseasonal variability $\tau$ accounts for several important atmospheric ENSO triggers, such as the westerly wind bursts (WWBs), the easterly wind bursts (EWBs), and the convective envelope of the Madden-Julian Oscillation (MJO) (Chen et al., 2015; Hu and Fedorov, 2016; Puy et al., 2016; Vecchi et al., 2006). Its strength is given by a state-dependent (multiplicative) noise that depends on the SSTa (Jin et al., 2007; Bianucci et al., 2018; Chen and Zhang, 2023), where a warmer SSTa leads to stronger wind burst activities. As for the decadal variability in the state variables through $I$, it stems from the observation that since 1870, through several detailed El Niño-type classification methods, the EP and CP El Niño events are alternatively prevalent every 10 to 20 years (Yu and Kim, 2013; Dieppois et al., 2021). This oscillation between EP-dominant and CP-dominant regimes indicates that the decadal variability plays an important role in the underlying dynamics, which is parameterized through a simple linear stochastic differential equation with multiplicative noise (10f), with no explicit dependence on the state variables in the faster time scales (Yang et al., 2021). They, together with additional small Gaussian white noise $\sigma_{u}\dot{W}_{u}$, $\sigma_{h}\dot{W}_{h}$, $\sigma_{C}\dot{W}_{C}$, $\sigma_{E}\dot{W}_{E}$, characterize the irregularity and multiscale features of the ENSO complexity (Timmermann et al., 2018; Fang and Xie, 2020).
@@
In the following, the non-dimensional form of the model in (10) is utilized to compute the responses. In the non-dimensional form, the six state variables have comparable maximum amplitudes. This means the strength of the components in the 6-dimensional column vector describing the eigenvector of the quadratic form can be used to intuitively tell the most sensitive direction. Nevertheless, the results for the evolution of the statistics with or without the perturbation are shown in the dimensional form that facilitates the physical explanations.